Q1. Given the upcoming NBA (professional basketball) draft, there are several players available with skills at different positions (e.g., point guard) and with or without college experience. The table below summarizes these data.

(Use the data below for this question only)


                                             College Experience (CE)   No College Experience (NCE)

Point Guard (PG)                                 20                                     39
Shooting Guard (SG)                           24                                     25
Center (C)                                            30                                     20
Small Forward (SF)                              25                                     13
Power Forward (PF)                             26 24

Find the following probability:

Q2.

Given the upcoming NBA (professional basketball) draft, there are several players available with skills at different positions (e.g., point guard) and with or without college experience. The table below summarizes these data.

(Use the data below for this question only)


                                             College Experience (CE)   No College Experience (NCE)

Point Guard (PG)                                 31                                     39
Shooting Guard (SG)                           13                                     37
Center (C)                                            25                                     16
Small Forward (SF)                              37                                     24
Power Forward (PF)                             39 37

Find the following probability:

p(C and NCE)

p(SG)

Q3.

Given the upcoming NBA (professional basketball) draft, there are several players available with skills at different positions (e.g., point guard) and with or without college experience. The table below summarizes these data.

(Use the data below for this question only)


                                             College Experience (CE)   No College Experience (NCE)

Point Guard (PG)                                 24                                     19
Shooting Guard (SG)                           19                                     25
Center (C)                                            10                                     25
Small Forward (SF)                              35                                     26
Power Forward (PF)                             37 24

Find the following probability:

p(CE)

Q4.

Given the upcoming NBA (professional basketball) draft, there are several players available with skills at different positions (e.g., point guard) and with or without college experience. The table below summarizes these data.

(Use the data below for this question only)


                                             College Experience (CE)   No College Experience (NCE)

Point Guard (PG)                                 39                                     16
Shooting Guard (SG)                           24                                     14
Center (C)                                            24                                     26
Small Forward (SF)                              29                                     19
Power Forward (PF)                             10 35

Find the following probability:

p(not C)

(If necessary, round answer to 4 decimal places)
Answer #1

Question 1:

The table is:

NCE College Experience(CE) No college Experience Total
Point Guard(PG) 20 39 59
Shooting Guard(SG) 24 25 49
Center C 30 20 50
Small Forward(SF) 25 13 38
Power Forward(PF) 26 24 50
total 125 121 246

p(not C) = 1-P(C) = 1-(50/246) = 1-0.2033 = 0.7967

The probability is 0.7967.

As per the Chegg answering guide, we have the option to answer only the first question in case of multiple questions. If you want to get the answers for the rest of the parts, please post the question in a new post.

Thank you! :)